OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 2x2 + 32x + 12 = 0

OpenStudy (anonymous):

First divide the whole equation by 2 to get the coefficient of x^2 to 1...

OpenStudy (anonymous):

First, move the untied number to the other side. (12) 2x2+32x=-12 Now divide by the term on x. (2) x2+16x=-6 Take half of the coefficient of the x term (16x) and square it. 8 ===> 64 Add this to both sides x2+16x+64=-6+64 Then...

OpenStudy (anonymous):

Take the left side and * factor * it. It's a perfect square trinomial. So you get (x+8)^2 And simplify the other side. You get 58. Our equation looks like (x+8)^2=58 :] I believe that's it. (srry I did the whole problem, I just wanted to clear up my memory on this)

OpenStudy (anonymous):

its fine thank you